OpenStudy (anonymous):

when the diagonal is 56 inches of a telivision,what is the length and breath

OpenStudy (dwelsher):

length =a^2 breath or width=b^2 b=(sq rt56-a^2) a=(sq rt 56-b^2) Is this correct??